# Efficient Nearest Neighbor Search Using Dynamic Programming

### v4

@@ V-E Other Types of Data
Figure 13: Average query time with varying sizes of the sampling box (arbitrary axis ratio, with volume ratio calculated as the cube of the axis ratio). Top: Uniform point cloud. Middle: Earth city data. Bottom: Dragon point cloud.
V-G Farthest point sampling
QuickFPS [22] is currently the fastest farthest point sampling strategy, achieving performance speedups of 230.5× over traditional CPU implementations, 43.4× over traditional GPU implementations, and 12.2× over previous state-of-the-art point cloud accelerators. It divides the point cloud into multiple buckets based on spatial location, and utilizes advanced techniques such as merged computation and implicit computation to accelerate the farthest point sampling process. Although the method proposed in this paper is not specifically designed for farthest point sampling, it also demonstrates impressive speed in this task. Using several traditional models as examples, with the farthest point sampling rate set at 25%, the sampling times of QuickFPS (CPU Version) and the method proposed in this paper are shown in Table IV. Moreover, compared to the complexity of QuickFPS, our proposed algorithm are both simpler and more intuitive.
@@ V-G Farthest point sampling
TABLE IV: Comparison of Sampling Times (s) between QuickFPS and Ours at a 25% Sampling Rate.
@@
QuickFPS 0.016 0.094 0.156 1.109 2.391 2.875 3.969
Ours 0.079 0.25 0.328 1.484 5.421 8.453 11.25
VI Desirable Properties
In light of its speed advantages, we believe that the current method harbors substantial potential for further advancements. In the following, we will delineate several pivotal aspects pertaining to its applications, aiming to elucidate the breadth and depth of its prospective utility in various domains.
VI-A Nearest Neighbor Among the First $k$ Points
One notable property of our algorithm is that, given any query point $q$ and integer $k$, we can efficiently identify the nearest point among $\{p_{i}\}_{i=1}^{k},k\leq n$. Specifically, the order of the given points is used to incrementally construct the Delaunay. As discussed in Section III, during the nearest-neighbor search process, when querying the nearest neighbor, we only need to compute $\Phi_{n}(q)$. However, in practice, $\forall~{}i\in[1,n]$, $\Phi_{i}(q)$ is already incorporated into our query process. When a specific $k$ is provided, as shown in Algorithm 2, our query process can terminate early. To the best of our knowledge, no existing algorithm is capable of achieving this.
The Density Peak Clustering algorithm [9] is a well-known method for clustering. One critical step in this algorithm is the calculation of $\sigma_{i}$, which is defined as:
$\delta_{i}=\min_{j:\rho_{j}>\rho_{i}}(d_{ij}),$ (9)
$\rho_{i}$ is defined as the density of point $p_{i}$, and $d_{ij}$ represents the distance between points $p_{i}$ and $p_{j}$. Generally, computing $\{\delta_{i}\}_{i=1}^{n}$ requires calculating the distance between every pair of points, which has a complexity of $O(n^{2})$. In contrast, our algorithm can efficiently solve this problem in $O(n\log n)$ time, assuming the data points are two-dimensional or three-dimensional. However, for higher-dimensional data, the time complexity of our algorithm increases due to the computationally expensive Delaunay triangulation process.
VI-B Farthest Point Sampling.
Farthest Point Sampling (FPS) is a widely used sampling algorithm because it ensures uniform point distribution, making it highly applicable across various tasks. For instance, FPS is used in the PointNet++ [35] framework for 3D point cloud deep learning to ensure uniform sampling and clustering of points, which contributes to building an effective receptive field. Although the initial purpose of our method was not specifically for farthest point sampling, it naturally lends itself to a simple and fast FPS strategy. Compared to existing methods, our proposed FPS strategy is more straightforward and easier to implement, offering an intuitive solution for efficient point sampling.
VI-C Various Primitives & Distance Metrics
Many existing nearest neighbor search algorithms are designed for specific types of primitives and distance metrics, which limits their generalizability across diverse scenarios. In contrast, our method, as described in Sections IIIIV-A, and IV-B, requires only the incremental construction of the Voronoi diagram for the primitives and a distance query function from any point in space to the primitives, making it straightforward to integrate into our nearest neighbor query framework. For instance, by supplying only the algorithm for incrementally constructing the Voronoi diagram of triangles under the Manhattan distance, along with a function to query the Manhattan distance between a spatial point and a triangle, our framework can efficiently determine the nearest primitive to a given point under the Manhattan distance. We believe this is a significant advantage of our algorithm. Of course, we acknowledge that the incremental construction of Voronoi diagrams for various primitives is quite challenging; however, as outlined below, we also provide alternative options.
VI-D Fast Preprocessing with Compromising Query Performance
In what follows, we aim to enhance the preprocessing efficiency at the expense of query performance. As elaborated in Section IV-A, during the construction of the Query Table, it is essential to identify the set $\omega_{m}$ upon inserting $p_{m}$, as illustrated in Algorithm 1. The set is defined by:
$\omega_{m}=\{p_{i}~{}|~{}p_{m}\vdash\text{Cell}(p_{i};\mathcal{V}_{m-1})\}.$ (10)
Indeed, when confronted with complex high-dimensional or intricate scenarios, computing $\omega_{m}$ becomes exceedingly challenging. Consequently, we opt to compute $\omega_{m}^{*}$, which merely needs to satisfy $\omega_{m}\subseteq\omega_{m}^{*}$, while still ensuring accurate nearest neighbor results. This approach opens up avenues for executing rapid nearest neighbor searches in high-dimensional spaces and diverse primitive and distance metric contexts. As an inevitable outcome, the computed set may harbor redundancies, thereby causing a marginal decrease in the speed of the nearest neighbor search.
Moreover, even if $\omega_{m}$ does not fall within $\omega_{m}^{*}$, it could furnish valuable insights conducive to the development of approximate nearest neighbor search algorithms.
VI-E Fast Query with Elaborated Preprocessing
Another encouraging finding is that the number of comparisons in our query process, as shown in Table I, can be further reduced, suggesting potential for additional algorithmic acceleration. Specifically, as previously outlined in our algorithm, when we focus on the time point $\Phi_{k}(q)=p_{k}$ during the query process, the current approach only utilizes the information that $q\in\text{Cell}(\Phi_{k}(q);\mathcal{V}_{k})$. However, we can infer that $q\in\cap_{\Phi_{i}(q)=p_{i},1<i\leq k}\text{Cell}(\Phi_{i-1}(q);\mathcal{V}_{i-1})~{}\cap~{}\text{Cell}(\Phi_{k}(q);\mathcal{V}_{k})$ from the query process.
As illustrated in Figure 14, (a) and (b) depict a simple incremental construction process, with the red dot representing the query point $q$. In this example, we have only utilized the information shown in (c), namely that $q\in\text{Cell}(p_{3};\mathcal{V}_{3})$. As shown in brown in (c). However, based on the query process, we can infer that $q\in\text{Cell}(p_{1};\mathcal{V}_{2})~{}\cap~{}\text{Cell}(p_{3};\mathcal{V}_{3})$, as is shown in (d). This enables us to design a more rigorous and precise evaluation criterion. Our method currently does not leverage this information during the query process, even though it can be obtained through complex preprocessing, which would, of course, involve substantial preprocessing overhead.
[…]
Figure 14: (a) and (b): A simple incremental insertion process, with the red dot representing the query point $q$. (c) Based on the query process of (a)(b), we know the information $q\in\text{Cell}(p_{3};\mathcal{V}_{3})$, as shown in brown. (d) In fact, the complete information we can obtain is $q\in\text{Cell}(p_{1};\mathcal{V}_{2})~{}\cap~{}\text{Cell}(p_{3};\mathcal{V}_{3})$, as shown in brown.
However, we also recognize that utilizing all of the aforementioned information would result in significantly increased preprocessing complexity. To balance this, we implemented a simplified approach that compromises between the algorithm described in this paper and the information mentioned in this section. In brief, in addition to $q\in\text{Cell}(p_{i};\mathcal{V}_{i})$, we incorporated $q\in\text{Cell}(\Phi_{i-1};\mathcal{V}_{i-1})~{}\cap~{}\text{Cell}(p_{i};\mathcal{V}_{i})$. This region is used as the fundamental unit for occupancy evaluation, rather than a whole Voronoi cell.
We were pleasantly surprised to find that, despite implementing only a simplified version, we achieved a significant reduction in the number of comparisons compared to Table I. As shown in Table V, $Ours^{1}_{*}$ and $Ours^{2}_{*}$ represent the number of comparisons without and with farthest point sampling, respectively, under the strategy outlined in this section. Remarkably, only around 100 pairwise distance comparisons are required to find the exact nearest neighbor for a query point among 1 million points. However, we must acknowledge that while the number of comparisons has decreased, the actual query time has increased. We speculate that this may be due to increased memory usage by the data structure during the query phase, which likely results in more cache misses, thereby reducing efficiency. Nevertheless, we believe this is a highly meaningful direction for further exploration.
TABLE V: The number of pairwise comparisons during the query process using different construction strategies.
Camel Bunny Dragon Kitten Armadillo Lucy Sponza
Vertices 28934 72911 100313 291023 726367 1018219 1313504
$\text{Ours}^{1}$ 171.783 166.726 209.516 182.719 234.742 276.549 418.285
$\text{Ours}^{2}$ 144.754 135.831 161.822 155.806 197.042 235.085 358.413
$\text{Ours}^{1}_{*}$ 93.218 97.49 108.545 107.708 127.760 134.123 124.615
$\text{Ours}^{2}_{*}$ 75.453 74.055 88.089 82.812 94.953 106.269 108.203
@@ VII Limitations and Future Work
The primary limitation of our algorithm is its high preprocessing cost, which restricts its applicability, particularly in large-scale scenarios or with high-dimensional datasets where efficiency is crucial. Therefore, in future work, we plan to couple the construction of our algorithm more closely with Delaunay computations, or alternatively, develop an approximate Delaunay construction to achieve faster preprocessing. These approaches aim to reduce the preprocessing burden and improve scalability for large-scale and high-dimensional applications.
Additionally, our method currently only supports nearest neighbor search. In future work, we aim to extend the algorithm to support k-nearest neighbor queries, which would enhance its versatility. Moreover, addressing all the desirable properties outlined in Section VI remains a significant long-term goal for us.
@@ VIII Conclusion
In this paper, we observe that traditional nearest neighbor search algorithms exhibit decreased efficiency when dealing with point clouds distributed over 2D manifolds in 3D space, or when query points are distant from the target points. In certain special cases, the complexity can degrade to $O(n)$. Inspired by the incremental Delaunay construction process, we propose a novel nearest neighbor search algorithm specifically designed for 2D manifold point cloud data. Our method achieves a speedup of 1-10 times compared to the KD Tree approach. We validated the efficiency advantages of our approach through extensive experiments. The high performance of our algorithm is also supported by evidence showing that it requires only about 100 pairwise distance comparisons to identify the point closest to the query point from a set of 1 million points.
Additionally, our algorithm demonstrates significant potential. For instance, it supports querying the closest point to $q$ among any of the first $k$ points. This capability facilitates the application of peak clustering algorithms to large-scale data scenarios. Moreover, we have implemented a farthest point sampling algorithm with $O(k\log n)$ complexity in 3D space. We believe that the algorithm presented in this paper holds significant potential for application and will make a substantial contribution to the community.

### v5

@@ VI-A Point-to-Mesh Distance Queries
Figure 17: Query time comparison for point-to-mesh distance computation on the ABC dataset.
TABLE V: Query time comparison ($\mu$s) of different methods for point-to-mesh distance computation on standard models. The highest and second-fastest query times are highlighted in bold and underlined, respectively
camel bunny armadillo dragon sponza kitten lucy
Faces 19510 69451 99976 249882 262267 268896 525814
Libigl 3.104 7.032 5.521 3.897 0.813 17.03 4.350
FCPW 0.933 1.984 1.617 1.061 0.213 4.537 1.338
[…]
VI-B Iterative Closest Point (ICP) Registration
Rigid registration, which finds an optimal rigid transformation to align a source point set with a target point set, is a fundamental problem in computer vision and many other areas. The Iterative Closest Point (ICP) algorithm [37] is one of the most important algorithms for rigid registration. ICP operates iteratively, where in each iteration, every point in the source point set requires querying its nearest neighbor in the target point set, which significantly impacts the algorithm’s efficiency.
In our experiments, we override the KD-tree implementation in PCL’s [30] ICP algorithm, substituting it with our method and nanoflann’s KD-tree implementation respectively. We evaluate three methods: the original ICP, ICP with our algorithm ($ICP_{ours}$), and ICP with nanoflann implementation ($ICP_{kdt}$) on 8 test scenarios from the 3DMatch dataset [38], performing registration on all adjacent frame pairs. The timing results are shown in Figure 18 (statistics include complete runtime including both preprocessing and querying), where the x-axis represents the sum of points in the source and target point clouds. Across all scenarios, our algorithm achieves faster registration in 82.4% of the cases, with speedups exceeding 10× compared to $ICP_{kdt}$ in the best cases. Moreover, the advantage becomes more pronounced as the number of queries increases.
[…]
Figure 18: ICP registration speed comparison on 8 3DMatch test cases. The horizontal axis shows the sum of target and source point cloud sizes. Our approach achieves faster performance on 82.4% of the data.
[…]
Figure 19: Four 2D clustering datasets from [39, 40], each containing 1M points.
VI-C Density Peak Clustering in 2D
Clustering is the task of grouping similar objects into clusters and represents a fundamental operation in data analysis and unsupervised machine learning. Clustering algorithms can be used to identify different types of tissues in medical imaging [41], analyze social networks, and identify weather regimes in climatology [42]. Among various clustering approaches, density-based clustering defines clusters as dense regions of points in the coordinate space and has received considerable attention. And Density peak clustering (DPC) [43] is a popular version of density-based clustering.
A critical step in the DPC algorithm involves calculating $\delta_{i}$, which is defined as:
$\delta_{i}=\min_{j:\rho_{j}>\rho_{i}}(d_{ij}),$ (8)
where $\rho_{i}$ represents the density of point $p_{i}$, and $d_{ij}$ denotes the distance between points $p_{i}$ and $p_{j}$. Generally, computing $\{\delta_{i}\}_{i=1}^{n}$ requires calculating distances between every pair of points, resulting in $O(n^{2})$ complexity. Numerous algorithms have been proposed to accelerate the computation of $\delta_{i}$ through heuristic approaches [44, 45].
One notable property of our algorithm is its ability to efficiently identify the nearest point among $\{p_{i}\}_{i=1}^{k},k\leq n$ for any query point $q$. As discussed in Section III, during the nearest-neighbor search process, when querying the nearest neighbor, we only need to compute $\Phi_{n}(q)$. However, in practice, $\forall~i\in[1,n]$, $\Phi_{i}(q)$ is already incorporated into our query process. When a specific $k$ is provided, as shown in Algorithm 2, our query process can terminate early. To the best of our knowledge, no existing algorithm achieves this capability.
Therefore, in 2D spaces, our method is particularly well-suited for accelerating density peak clustering algorithms. We simply sort the point cloud by density and use this ordering for incremental preprocessing in our algorithm. In our experiments, we replace the $\delta$ computation process in Ex-DPC++ [44], the fastest known density peak clustering algorithm as we known, with our method. We evaluate performance on four classical scenarios [39, 40], each containing 100,000 points, as shown in Figure 19, and 1000 randomly generated 2D clustering scenarios. Table VI and Figure 20 present the comparative results. Without parallelization, our clustering achieves over 10× speedup compared to the original Ex-DPC++. Even with parallelization enabled (while our $\delta$ computation remains serial), our method maintains 1-3× speed advantages. It should be noted that due to the complexity of high-dimensional Delaunay construction, our strategy-accelerated density peak clustering algorithm is currently suitable only for low-dimensional spaces (2D/3D), with high-dimensional clustering requiring further investigation.
Beyond the speed advantages, our method offers stronger theoretical foundations compared to existing heuristic-based and parallelized algorithms. Additionally, our capability to query the nearest point among the first $k$ points may find applications in various practical scenarios:
  1. 1.
    Finding the most similar weather conditions to current patterns that occurred before a specific date for climate analysis.
  2. 2.
    Identifying the closest historical stock price pattern to current market conditions, restricted to periods before specific regulatory changes.
  3. 3.
    In mobile location services, users seek restaurants that are both “affordable and nearby.” The system needs to identify the nearest restaurant among all lower-priced establishments..
  4. 4.
    Finding “highly qualified and role-appropriate” benchmark employees for job candidates during recruitment processes.
  5. 5.
    Journal editors need to identify reviewers who are both “academically distinguished and research-field relevant” for submitted manuscripts.
[…]
Figure 20: Clustering speed comparison on 1000 synthetic clustering data examples, without (up) and with (bottom) parallelization. Our method demonstrates consistent performance advantages in both scenarios.
TABLE VI: Clustering speed comparison (in seconds) on four clustering data examples, before and after OpenMP parallelization. Our method demonstrates consistent speed advantages across all examples.
syn-2d Birch1 Birch2 Birch3
Ours 0.715 0.432 0.577 0.796
Ex-DPC++ 4.974 5.723 4.961 5.386
Ours (omp) 0.559 0.298 0.407 0.610
Ex-DPC++ (omp) 0.927 0.953 0.1008 0.960
VI-D Point-to-Segments Distance Query
In 2D scenarios, computing the nearest distance from a point to a collection of line segments represents a fundamental task in computer graphics, with widespread applications in path planning, navigation, and collision detection. Traditional approaches for nearest segment queries typically employ hierarchical bounding volume strategies, such as bounding volume hierarchies (BVH) or R-trees, to efficiently organize and traverse the line segment collection.
Although this paper primarily focuses on computing nearest distance queries from points to point sets on 2D manifold surfaces in 3D space, our nearest neighbor search framework can be extended to different primitives, dimensions, distance metrics, and even solution spaces, provided that incremental Delaunay construction algorithms exist for the specific scenario and a distance query function from any point in the computational space to the primitives is available. Leveraging CGAL’s integrated 2D segment Delaunay construction algorithm [46], our nearest neighbor search approach can be adapted for point-to-disjoint-segments nearest distance queries in 2D.
In our experiments, we randomly generate 50,000 non-intersecting line segments within the domain $[-5,5]^{2}$, with segment lengths randomly distributed in the range $[0,\delta]$. By controlling the parameter $\delta$, we obtain multiple test scenarios with varying geometric complexity. We randomly generate 1 million query points within $[-5,5]^{2}$ and compare against FCPW and R*Tree (from Boost). Figure 21 presents the comparative results, where the x-axis represents the percentage of $\delta$ relative to the diagonal length of the generation space ($5\sqrt{2}$), and the y-axis shows the average query time per operation. As observed, both R*Tree and BVH-based methods exhibit declining query efficiency as the maximum segment length increases. In contrast, our method maintains consistent query efficiency across scenarios with different maximum segment lengths while demonstrating superior performance compared to existing approaches. Preprocessing times are shown in Figure S2 of the supplementary material.
[…]
Figure 21: Point-to-segments distance query performance comparison with varying maximum segment lengths.
VI-E Farthest Point Sampling as a Natural Extension
Although our method was not specifically designed for farthest point sampling, our incremental Voronoi construction framework naturally extends to support this functionality. The algorithm can readily identify farthest points during the incremental construction process, making it a straightforward derivation of our core approach.
While our naturally derived implementation is slower than existing specialized farthest point sampling algorithms [47], it demonstrates the versatility of our framework and provides a simple solution that emerges directly from our Voronoi construction methodology. The comparative performance results are presented in Table VII.
@@ VI-E Farthest Point Sampling as a Natural Extension
TABLE VII: Farthest point sampling time comparison between QuickFPS and our method at 25% sampling rate on standard models.
@@
QuickFPS 0.023 0.064 0.087 0.265 0.865 1.401 2.013
Ours 0.035 0.102 0.142 0.610 2.361 3.428 4.501
@@ VII Limitations and Future Work
The primary limitation of our algorithm is its high preprocessing cost and the difficulty of parallelizing the preprocessing phase, which restricts its applicability, particularly in large-scale scenarios or with high-dimensional datasets where efficiency is crucial. Therefore, in future work, we plan to couple the construction of our algorithm more closely with Delaunay computations, or alternatively, develop an approximate Delaunay construction to achieve faster preprocessing. Additionally, we will explore the potential for accelerating the preprocessing phase through parallelization strategies. These objectives aim to reduce the preprocessing burden and improve scalability for large-scale and high-dimensional applications.
Additionally, our method currently only supports nearest neighbor search. In future work, we aim to extend the algorithm to support k-nearest neighbor queries, which would enhance its versatility. Furthermore, our approach is primarily effective in low-dimensional spaces (2D/3D), as high-dimensional Delaunay triangulation becomes computationally expensive with prohibitive preprocessing costs. Exploring approximate Delaunay construction techniques for high-dimensional spaces represents another promising direction that could potentially overcome this dimensional constraint while maintaining reasonable preprocessing costs.
@@ VIII Conclusion
In this paper, we observe that traditional nearest neighbor search algorithms exhibit decreased efficiency when dealing with point clouds distributed over 2D manifolds in 3D space, or when query points are distant from the target points. In certain special cases, the complexity can degrade to $O(n)$. Inspired by the incremental Delaunay construction process, we propose a novel nearest neighbor search algorithm specifically designed for 2D manifold point cloud data. Our method achieves 1-10× speedup compared to KD-Tree and R*-Tree approaches, with efficiency advantages validated through extensive experiments.
Additionally, our algorithm demonstrates significant practical potential. We validated the effectiveness and versatility of our approach through five diverse applications: point-to-mesh distance queries, ICP registration, 2D density peak clustering, point-to-segments distance queries, and farthest point sampling. We believe that the algorithm presented in this paper offers substantial practical value and theoretical contributions that will benefit the broader research community.